# CoAM: Corpus of All-Type Multiword Expressions

### v1

@@ 4.6 Review
Type PoS of Head Description Example
Verb VERB Verbal MWEs. stand for, pick up, break a leg
Mod/ Conn ADJ, ADV, ADP, CCONJ, SCONJ Adjectival, adverbial, or connective MWEs. under the weather, of course, in spite of
Clause VERB, AUX MWEs containing (1) a verb or auxiliary verb and (2) its nominal subject, e.g., proverbs, phatic expressions. you’re welcome, the early bird gets the worm, when it comes to
Other Any MWEs whose head is not contained in them or whose PoS is none of the PoS above. and so on
@@
Would recomend giving this a a try.
@@ 4.7 Consistency Check
We found 118 instances of inconsistencies like this in the random sample of DiMSUM and 147 in CoAM,888 When counting inconsistencies, we always consider positive labels correct, and negative labels inconsistent. reaffirming the difficulty involved in producing consistent MWE annotations. However, we then added the missing labels to all MWEs found in this way, eliminating these inconsistencies from the final CoAM data.
@@
News 0360 09,328 230 5.5 33.5 47.8 16.5 0.0 2.2
Commentary 0357 09,310 272 7.0 29.8 30.5 29.4 1.1 9.2
TED 0299 06,592 212 7.2 33.0 37.3 21.2 3.8 4.7
UD 0285 05,001 160 7.2 - - - - -
Train 0780 16,817 489 6.7 33.4 35.9 22.8 1.5 6.4
Test 0521 13,414 385 6.5 30.6 40.0 22.9 1.6 4.9
Total 1,301 30,231 874 6.6 31.9 38.1 22.8 1.5 5.6
@@ 4.8 MWE Type Tagging
To enable fine-grained error analysis using CoAM, we automatically tagged all test-set MWEs, and training-set MWEs except those from UD with MWE types. Inspired by the classification by Schneider et al. (2014), we group MWEs into five types: Noun, Verb, Modifier/Connective (Mod/Conn), Clause, and Other. The details are described in Table 2. We tag MWEs in each sentence $s$ through the following automatic operations after dependency parsing using SpaCy. For each MWE $m$ in $s$, we look for the head word, namely, try to find $w$ that has all the other words in $m$ as its descendants. If $m$ has such $w$, we determine its type based on the PoS of $w$; for example, we tag $m$ as Noun when $w$ is a noun. If $m$ does not have such $w$, we tag $m$ as Other. We did not tag the sentences in UD because none of them are contained in the test set.
@@ 4.9 Statistics
Table 4 shows dataset statistics. CoAM has more than 1.3K sentences. The MWE density is 6–7% in both training and test sets. This proportion is much lower than that of DiMSUM Schneider et al. (2016)—13% in the grand total—most likely because they included proper noun phrases in MWEs. The right five columns show that Verb and Noun are the two most frequent MWE types across data sources. They also demonstrate that our type tagging approach successfully assigned a specific (non-Other) MWE type to almost 95% of the MWEs in CoAM.
We also report the ratio of unseen MWEs, because they are of interest to the MWE research community, with their identification being a major challenge Ramisch et al. (2020). An MWE in the test set is considered unseen if the multi-set of lemmas of its constituents was never annotated in the training set. Among 385 MWEs in our test set, 64.2% are unseen.
@@ 5.1 MWEasWSD
MWEasWSD (MaW) is a system that uses (1) an MWE lexicon—WordNet Miller (1995)—and a rule-based pipeline to identify MWE candidates and (2) a trainable model to filter MWE candidates Tanner and Hoffman (2023). It achieved state-of-the-art performance on the DiMSUM dataset. They published four of their filtering models, among which we use the bi-encoder (BiEnc) and DCA poly-encoder (DCA). Both models are based on BERT Devlin et al. (2019), specifically bert-base-uncased. They have been trained with SemCor Miller et al. (1993), and we further fine-tune each model with the CoAM training set. We also run MaW with the rule-based pipeline only, i.e., without a filtering model.
@@ 5.2 LLM Fine-Tuning
Recent studies have shown the effectiveness of LLM fine-tuning for a wide range of tasks, such as NER Zhou et al. (2024) and grammatical error correction Kaneko and Okazaki (2023). These works achieved high performance by providing LLMs with prompts that included instructions on the task.
@@ 6.1 Setup
Recall by MWE Type
F1 P R Noun Verb Mod/Conn Clause
(118) (154) (88) (6)
FT Llama-8B 29.4±2.1 $\mathbf{82.6}$±1.2 17.9±1.6 5.9±0.8 26.4±1.5 24.2±3.5 0.0±0.0
Llama-70B 38.4±4.8 74.5±2.6 26.1±4.6 19.2±4.2 35.7±5.7 25.4±5.6 0.0±0.0
Qwen-7B 45.2±0.6 63.2±0.9 35.2±0.8 26.0±2.0 47.2±1.0 31.8±2.0 0.0±0.0
Qwen-72B $\mathbf{55.5}$±0.5 61.5±2.6 $\mathbf{50.7}$±2.5 $\mathbf{44.4}$±3.2 $\mathbf{60.6}$±1.0 $\mathbf{50.4}$±5.1 $\mathbf{22.2}$±9.6
MaW Rule 32.7 28.3 38.7 37.3 40.9 47.7 0.0
Rule+BiEnc 41.6±0.1 48.6±0.3 36.5±0.3 32.2±0.0 41.1±0.7 44.3±0.0 0.0±0.0
Rule+DCA 42.0±0.1 48.4±0.6 37.1±0.3 33.3±0.5 40.9±0.0 45.8±0.7 0.0±0.0
Recall by Seen/Unseen Recall by In WN or Not
Seen Unseen True False
(138) (247) (163) (222)
FT Llama-8B 37.2±3.4 7.2±0.6 29.2±1.9 9.6±1.7
Llama-70B 35.0±7.3 21.1±3.2 46.8±6.5 10.8±3.2
Qwen-7B 44.4±0.8 30.0±0.8 50.1±1.4 24.2±0.5
Qwen-72B $\mathbf{58.2}$±5.8 $\mathbf{46.6}$±0.7 63.6±3.4 $\mathbf{41.3}$±1.9
MaW Rule 47.8 33.6 $\mathbf{91.4}$ 0.0
Rule+BiEnc 44.2±0.0 32.1±0.5 86.1±0.7 0.0±0.0
Rule+DCA 45.4±0.8 32.4±0.0 87.5±0.7 0.0±0.0
@@
LLM Fine-Tuning
We use four instruction-tuned LLMs that are available on Hugging Face Hub: Llama-3.1-8B-Instruct, Llama-3.1-70B-Instruct Dubey et al. (2024), Qwen-2.5-7B-Instruct, and Qwen-2.5-72B-Instruct Qwen Team (2024). We hereafter abbreviate them, e.g., to Llama-8B, omitting the versions. For efficient training and inference, we use QLoRA Dettmers et al. (2023), performing 4-bit NormalFloat quantization and double quantization. Other hyperparameters are described in Section B.1. The computational budgets are in Appendix C.
To find a performant input-output format, we conduct preliminary experiments with three formats, adjusting the prompt for each format. These experiments are detailed in Section B.1. We find tsv_to_tsv (see Table 6) is the only format the models can comply with.
All prompts contain a placeholder for a definition of MWEs to clarify what sequences should be marked. We report scores using what we refer to as the long definition in Table 5; for scores using a shorter definition, see the ablation study in Section B.1.
@@
TP fire up The allegations have fired up the opposition, … Verb MWE in WordNet
TP at least … since at least the 1950s. Mod/Conn MWE in WordNet
FN real estate …park their toxic real estate assets … Noun MWE not in WordNet
FN you know You know, it’s very old … Clause MWE not in WordNet
@@ 6.2 Results and Analysis
In Table 5a, the left three columns show the overall scores. Fine-tuned Qwen-72B achieves the best F1, surpassing the highest F1 by Rule+DCA of MaW. Moreover, fine-tuned Qwen-72B outperforms all MaW systems in precision and recall. This indicates the effectiveness of fine-tuning LLMs for MWEI—particularly LLMs with a large number of parameters. One explanation for this could be that knowledge about MWEs was acquired by LLMs through pre-training, and we can harness this knowledge for MWEI by fine-tuning.
Meanwhile, all systems except for the rule-based pipeline suffer from low recall. Even the best system, fine-tuned Qwen-72B, achieves a recall of only 50.7%, failing to identify almost half of the gold MWEs. This leads us to conduct further analysis.
@@
The right columns in Table 5a compare recall by MWE type, showing that Clause and Noun MWEs are more difficult to identify than Mod/Conn or Verb MWEs across models. In Table 5b, the left columns show that seen MWEs are easier to identify than unseen ones across models/systems. This result resembles that of the PARSEME shared task 1.2 Ramisch et al. (2020), demonstrating that unseen MWEs remain challenging in MWEI.
The right columns in Table 5b reveal that MWEs not contained in WordNet are much more difficult to identify than those in WordNet. It is natural that MaW systems cannot identify the MWEs not in WordNet, as their rule-based pipeline cannot detect these candidates, but interestingly we find that fine-tuned LLMs also struggle to identify MWEs not in WordNet. For Qwen-72B, the F1 score gap between MWEs contained and not contained in WordNet is about 22 points, much larger than the gap between seen/unseen MWEs, about 12 points. We hypothesize that this is caused by MWEs in WordNet being more widely recognized as idioms and that this recognition is reflected in the training data of the LLMs, enhancing their ability to identify these MWEs.
Table 7 shows examples of correctly identified MWEs and missed MWEs by fine-tuned Qwen-72B. 79.0% of Verb MWEs in WordNet like fire up are correctly identified by the model. Meanwhile, the Noun MWE real estate was not identified. The aforementioned hypothesis could explain this, as real estate is not in WordNet. Accurately identifying such multi-word entities or multi-word terms Savary et al. (2019) could be a challenge for future studies.
@@
Given the high performance of fine-tuned Qwen-72B, we investigate how much fine-tuning (FT) contributes to the performance by comparing FT to zero-shot learning (ZSL) and few-shot learning (FSL). In ZSL, we provide the models with the same prompt as FT. In FSL, we sample 5 pairs of input (sentence) and output (gold MWE set) from the training set, convert them into the tsv_to_tsv format, and include them in the prompt as examples. In the sampling process, we repeat random sampling until at least two of the example sentences contain one or more MWEs so that the models have a sufficient number of examples to learn the task and format.
Table 8 shows the results. FT greatly outperforms ZSL and FSL, demonstrating the effectiveness of FT on the CoAM training set.
@@
FT 38.4±4.8 55.5±0.5
FSL 4.3±0.9 14.1±0.3
ZSL 6.9 2.8
Table 8: Ablation results by MWEI method and model, as mean percentage F1 scores of three runs (for FT the randomness arises in training, and for FSL in exemplar selection). $\pm$ denotes standard deviation.
Relative Strengths of LLMs vs MWEasWSD
As shown in Table 5a, fine-tuned Qwen-72B outperforms the best MaW system, Rule+DCA, in all of F1, precision, and recall, by over 10 points. However, it does not mean fine-tuning should always be preferred. First, MaW’s encoder model (bert-base-uncased) has only 110M parameters, taking approximately $2\times 110\textrm{M}\times 4=880\textrm{M}$ bytes in total ($2$ is the number of encoders). 4-bit quantized Qwen-72B takes $72,000\textrm{M}\times 0.5=36,000\textrm{M}$ bytes, taking roughly 40 times more space than MaW. On the other hand, the performance of 4-bit quantized Qwen-7B, which takes roughly 4 times more space, is comparable with Rule+DCA, suggesting that MaW is more memory-efficient.
Furthermore, the bottleneck of MaW is identified to be the size of the lexicon Tanner and Hoffman (2023). Future work could significantly improve the system’s performance simply by using a lexicon containing more MWEs than WordNet.
@@ 7 Conclusion
In this paper, we constructed CoAM, a high-quality dataset of 1.3K sentences for MWEI covering all types of MWEs. Using a combination of human review and automated consistency checking, we addressed consistency issues that have been a problem for previous MWE datasets, enabling more accurate evaluation results for future work in MWEI. We used CoAM to evaluate two MWEI approaches: MaW and LLM fine-tuning. Our largest fine-tuned LLM performed the best, outperforming the current state-of-the-art, but all systems suffered from low recall. Consequently, we argue that MWEI overall remains a challenging task.
8 Limitations
Despite our best efforts to craft clear and unambiguous annotation guidelines, our initial annotations on CoAM suffered from very low inter-annotator agreement, suggesting possible inconsistencies. We have done our best to address this problem through the combination of human review (Section 4.6) and consistency checking (Section 4.7), but we acknowledge that in an ideal world initial inter-annotator agreement would have been higher.
Additionally, our annotation interface generator, CAIGen suffers from a few drawbacks. For example, it requires annotators to change rows when they have marked some MWEs and are marking more. Because annotators can forget to change rows, which happened in the construction of CoAM, researchers should urge them to confirm their annotations are as intended after finishing each sentence. Taking this into consideration, we release CAIGen’s source codes so that potential users can judge for themselves if its benefits outweigh its drawbacks.
9 Ethical Considerations
All sources of CoAM are permitted at minimum for use in research, as described in Section A.1. CoAM itself will be released under the condition that the users do not publish the data on the open web to prevent its leakage to the training data of future models. All annotators of CoAM consented that their annotations be made publicly accessible. Throughout the dataset construction processes involving authors, we found no contents harmful to the environment, specific groups of people, privacy, and so on.
10 Acknowledgments
To paraphrase or polish authors’ original content, we used an AI writing assistant tool, ChatGPT.999 https://chatgpt.com Financially, this work was supported by JST SPRING Grant Number JPMJSP2140.

### v2

@@ 4.6 Review
Would recomend giving this a try.
@@ 4.7 Consistency Check
We found 118 instances of inconsistencies like this in the random sample of DiMSUM and 147 in CoAM,999 When counting inconsistencies, we always consider positive labels correct, and negative labels inconsistent. reaffirming the difficulty involved in producing consistent MWE annotations. However, we then added the missing labels to all MWEs found in this way, eliminating these inconsistencies from the final CoAM data.
@@
News 360 9,328 230 5.5 32.2 44.3 23.0 0.0 0.4
Commentary 357 9,310 269 7.0 30.5 29.7 33.1 1.5 5.2
TED 299 6,592 210 7.1 35.2 38.6 21.4 3.8 1.0
UD 285 5,001 158 7.2 42.4 39.2 12.7 1.3 4.4
Train 780 16,817 486 6.7 36.2 38.3 19.8 1.4 4.3
Test 521 13,414 381 6.4 31.8 36.5 29.1 1.8 0.8
Total 1,301 30,231 867 6.6 34.3 37.5 23.9 1.6 2.8
@@ 4.8 MWE Type Tagging
To enable fine-grained error analysis, we automatically tagged all the MWEs in CoAM with MWE types. We then manually corrected the tags in the test set, as described below, to enable a precise analysis of the results by MWE types.
Inspired by the classification by Schneider et al. (2014b), we group MWEs into five types: Noun, Verb, Modifier/Connective (Mod/Conn),101010 We include adjectival, adverbial, adpositional, and connective MWEs in the category of Mod/Conn to create a category with a sufficient data size. Clause, and Other. The details are described in Table 2. We automatically tag MWE types in each sentence based on the dependency structure. Dependency parses were provided in the data source for UD and obtained using spaCy for News, Commentary, and TED. For each MWE $\bm{m}=(w_{1},\ldots,w_{|\bm{m}|})$, we look for its syntactic head, namely, try to find $w^{*}$ that has all the other words in $\bm{m}$ as its descendants. If $\bm{m}$ has such $w^{*}$, we determine its type based on the PoS of $w^{*}$; for example, we tag $\bm{m}$ as Verb when $w^{*}$ is a verb. If $\bm{m}$ does not have such $w^{*}$, we tag $\bm{m}$ as Other. For the details, see the algorithm in Section A.3.
We then manually checked all 384 MWEs in the test set and corrected the errors in them. Two authors who are native English speakers conducted this, identifying and correcting 30 errors. The error rate of automatic tagging was therefore 7.8%. The most common errors were Mod/Conn automatically tagged as Verb (e.g., according to) and Other (e.g., not only) constituting 3.3% and 3.1% of the test data, respectively. We expect this to be representative of the errors in the type tags automatically assigned to the training set. Note, however, that we used only the manually corrected tags on the test set for result analysis in Section 6.2.
@@ 4.9 Statistics
Table 4 shows dataset statistics. CoAM has more than 1.3K sentences. The MWE density is 6–7% in both training and test sets. This proportion is much lower than that of DiMSUM Schneider et al. (2016)—13% overall—most likely because DiMSUM includes proper noun phrases in MWEs. Among the five MWE types, Verb and Noun are the most frequent ones across data sources. See Section A.4 for additional analysis regarding MWE continuity.
@@ 5.1 MWEasWSD
MWEasWSD (MaW) is an approach that uses (1) an MWE lexicon—WordNet Miller (1995)—and a rule-based pipeline to identify MWE candidates and (2) a trainable model to filter MWE candidates Tanner and Hoffman (2023). It achieved state-of-the-art performance on the DiMSUM dataset. They published four of their filtering models, among which we use the bi-encoder (BiEnc) and DCA poly-encoder (DCA). Both models are based on BERT Devlin et al. (2019), specifically bert-base-uncased. They have been trained with SemCor Miller et al. (1993), and we further fine-tune each model with the CoAM training set. We also run MaW with the rule-based pipeline only, i.e., without a filtering model.
@@ 5.2 LLM Fine-Tuning
Recent studies have shown the effectiveness of LLM fine-tuning for a wide range of tasks, such as NER Zhou et al. (2024) and grammatical error correction Kaneko and Okazaki (2023). Framing their task as sequence transduction, they achieved high performance by providing LLMs with prompts that included instructions on the task.
@@ 6.1 Setup
LLM Fine-Tuning
We use four instruction-tuned LLMs that are available on Hugging Face Hub: Llama-3.1-8B-Instruct, Llama-3.1-70B-Instruct Dubey et al. (2024), Qwen-2.5-7B-Instruct, and Qwen-2.5-72B-Instruct Qwen Team (2024). We abbreviate them, e.g., to Llama-8B, omitting the versions. For efficient training and inference, we use QLoRA Dettmers et al. (2023), performing 4-bit NormalFloat quantization and double quantization. Other hyperparameters are described in Section B.1. The computational budgets are in Appendix C.
To find a performant input-output format, we conduct preliminary experiments with three formats, adjusting the prompt for each format. These experiments are detailed in Section B.1. We find tsv_to_tsv (see Table 5) is the only format the models can comply with and thus adopt it.
All prompts contain a placeholder for a definition of MWEs to clarify what sequences should be marked. We report scores using what we refer to as the long definition in Table 6. For scores using a shorter definition, see the ablation study in Section B.1.
@@
Recall by MWE Type
F1 P R Noun Verb Mod/Conn Clause
(121) (139) (111) (7)
FT Llama-8B 24.9±3.7 $\mathbf{92.0}$±2.0 14.4±2.4 3.3±0.0 19.4±2.9 21.6±4.8 0.0±0.0
Llama-70B 38.3±5.2 69.0±1.7 26.8±5.1 19.0±5.2 39.6±7.5 21.6±2.7 0.0±0.0
Qwen-7B 48.1±1.0 60.9±0.8 39.7±1.0 28.7±1.3 52.0±0.8 39.6±0.9 0.0±0.0
Qwen-72B $\mathbf{57.8}$±1.8 63.8±2.2 $\mathbf{52.8}$±2.1 $\mathbf{42.1}$±4.4 $\mathbf{59.7}$±1.4 $\mathbf{57.7}$±1.8 $\mathbf{28.6}$±0.0
MaW Rule 32.4 27.9 38.6 $\mathbf{33.1}$ $\mathbf{45.3}$ $\mathbf{39.6}$ 0.0
Rule+BiEnc 41.4±0.3 47.9±0.8 36.5±0.3 28.9±0.0 45.3±0.0 36.9±0.9 0.0±0.0
Rule+DCA 41.9±0.2 49.0±0.5 36.7±0.5 29.8±0.0 45.3±0.7 36.6±1.0 0.0±0.0
Recall by Seen/Unseen Recall by Continuity Recall by In WN or Not
Seen Unseen Continuous Discontinuous True False
(141) (240) (338) (43) (161) (220)
FT Llama-8B 29.6±5.0 5.6±0.9 16.0±2.7 2.3±0.0 23.4±3.2 7.9±1.8
Llama-70B 37.1±4.5 20.7±5.4 29.0±4.9 9.3±6.2 43.9±6.3 14.2±4.1
Qwen-7B 47.3±0.4 35.3±1.3 43.9±1.1 7.0±0.0 53.8±0.9 29.4±1.1
Qwen-72B $\mathbf{61.9}$±0.8 $\mathbf{47.4}$±3.8 $\mathbf{57.3}$±1.8 $\mathbf{17.1}$±4.8 62.7±2.8 $\mathbf{45.5}$±2.8
MaW Rule 46.8 33.8 41.7 14.0 $\mathbf{91.3}$ 0.0
Rule+BiEnc 45.2±0.4 31.4±0.2 39.6±0.3 11.6±0.0 86.3±0.6 0.0±0.0
Rule+DCA 44.0±1.2 32.4±0.2 40.3±0.5 7.8±1.3 86.7±1.3 0.0±0.0
@@
TP fire up The allegations have fired up the opposition, … Verb in WordNet
TP at least … since at least the 1950s. Mod/Conn in WordNet
FN in …hands …concentration of power in his own hands. Mod/Conn in WordNet
FN real estate …park their toxic real estate assets … Noun not in WordNet
FN you know You know, it’s very old … Clause not in WordNet
@@ 6.2 Results and Analysis
In Table 6a, the left three columns show the overall scores. Fine-tuned Qwen-72B achieves the best F1, surpassing the highest F1 by Rule+DCA of MaW. Additionally, fine-tuned Qwen-72B outperforms all MaW systems in precision and recall. This indicates the effectiveness of fine-tuning LLMs for MWEI—particularly LLMs with a large number of parameters. One explanation for this could be that knowledge about MWEs was acquired by LLMs through pre-training, and we can harness this knowledge for MWEI by fine-tuning.
Meanwhile, all systems, except for the rule-based one, suffer from low recall. Even the best system, fine-tuned Qwen-72B, achieves a recall of only 52.8%, missing almost half of the gold MWEs.
@@
Given the low recall across all systems, we conduct further analysis to determine which MWEs the models particularly struggle to identify.
Recall by MWE Type in Table 6a shows that Clause and Noun MWEs tend to be more difficult to identify than Mod/Conn or Verb MWEs across models.
In Table 6b, the first two columns analyze how much recall changes depending on whether the MWE is seen or unseen, where an MWE in the test set is considered unseen if the multi-set of lemmas of its constituents was never annotated in the training set. We find that seen MWEs are easier to identify than unseen ones across models and systems, which resembles the results of the PARSEME shared task 1.2 Ramisch et al. (2020).
Recall by Continuity in Table 6b shows that discontinuous MWEs are, across models and systems, far more difficult to identify than continuous ones, indicating that discontinuous MWEs remain a key challenge and require further attention.
Recall by In WN or Not in Table 6b reveals that MWEs not contained in WordNet are much more difficult to identify than those in WordNet. It is natural that MaW systems cannot identify the MWEs not in WordNet, as their rule-based pipeline cannot detect these candidates, but interestingly we find that fine-tuned LLMs also struggle to identify MWEs not in WordNet. We hypothesize that this is caused by MWEs in WordNet being more widely recognized as MWEs or idioms and that this recognition is reflected in the training data of the LLMs, enhancing their ability to identify these MWEs.
Table 7 shows examples of correctly identified MWEs and missed MWEs by fine-tuned Qwen-72B. Verb MWEs in WordNet like fire up are relatively easy to identify, with a recall of 79.0% achieved by the model. Meanwhile, the discontinuous MWE, in …hands, was not correctly identified. Additionally, the Noun MWE real estate was not identified; the aforementioned hypothesis could explain this, as real estate is not in WordNet. Accurately identifying those MWEs could be a challenge for future studies.
@@
Given the relatively high performance of fine-tuned Qwen-72B, we investigate how much fine-tuning (FT) contributes to the performance by comparing FT to zero-shot learning (ZSL) and few-shot learning (FSL). In ZSL, we provide the models with the same prompt as FT. In FSL, we sample 5 pairs of input (sentence) and output (gold MWE set) from the training set, convert them into the tsv_to_tsv format, and include them in the prompt as examples. In the sampling process, we ensure that the models have a sufficient number of examples to learn the task and format by repeating random sampling until at least two of the example sentences contain one or more MWEs.
Table 8 shows the results. FT greatly outperforms ZSL and FSL, demonstrating the effectiveness of fine-tuning on the CoAM training set.
@@
FT 38.3±5.2 57.8±1.8
FSL 6.1±0.0 12.9±0.0
ZSL 3.9 1.4
Table 8: Ablation results by MWEI method and model, as mean percentage F1 scores of three runs (for FT the randomness arises from training, and for FSL from example selection). $\pm$ denotes standard deviation.
Comparison considering computational efficiency
As shown in Table 6a, fine-tuned Qwen-72B outperforms the best MaW system in all of the F1 score, precision, and recall. However, the LLM takes substantial memory and compute, and MaW is more efficient in this regard. MaW’s encoder model (bert-base-uncased) has only 110M parameters, taking approximately $2\times 110\textrm{M}\times 4=880\textrm{M}$ bytes in total ($2$ is the number of encoders). 4-bit quantized Qwen-72B takes $72,000\textrm{M}\times 0.5=36,000\textrm{M}$ bytes, taking roughly 40 times more space than MaW. On the other hand, the performance of 4-bit quantized Qwen-7B, which takes roughly 4 times more space, is comparable with Rule+DCA.
@@ 7 Conclusion
In this paper, we constructed CoAM, a high-quality dataset of 1.3K sentences for MWEI covering all types of MWEs. Using a combination of human review and automated consistency checking, we addressed consistency issues that have been a problem for previous MWE datasets, enabling more accurate evaluation results for future work in MWEI. We used CoAM to evaluate two MWEI approaches: MaW and LLM fine-tuning. Our largest fine-tuned LLM performed the best, but all systems suffered from low recall. Consequently, we argue that MWEI overall remains a challenging task.
Limitations
Dataset Size
Because the objective of this work was to maximize dataset quality, we invested heavily in efforts to improve data quality, employing multiple annotators per sentence, manual review, and consistency checking. This focus on quality over quantity resulted in a slightly smaller dataset size (1.3K sentences) compared to some previous works like DiMSUM and PARSEME. However, the evaluation set of CoAM has many more sentences than the gold data of ID10M Tedeschi et al. (2022), and is large enough for our purpose of reliably evaluating MWEI systems.
CAIGen
We observed a usability issue with our annotation interface, CAIGen. It requires annotators to use separate table rows for each annotated MWE. Because annotators may forget to use separate rows, which happened in the construction of CoAM, researchers should urge them to confirm their annotations are as intended after finishing each sentence.
MWEI Approaches
In this study, we experimented with an existing approach, MaW, as well as LLM fine-tuning. Future work could explore other previously proposed methods, such as BERT-based tagging with BIO-like labels, as introduced by Taslimipoor et al. (2020).
Ethical Considerations
All sources of CoAM are permitted at minimum for use in research, as described in Section A.1. CoAM itself will be released under the condition that the users do not publish the data on the open web to prevent its leakage to the training data of future models. All annotators of CoAM consented to the publication of their annotations. Throughout the dataset construction processes, we found no content harmful to the environment, specific groups of people, or privacy.
Acknowledgments
We are grateful for the insightful comments by Yuki Arase, Shohei Higashiyama, and anonymous reviewers. To polish the content of this paper, we used ChatGPT111111 https://chatgpt.com as a writing assistant tool. Financially, this work was supported by JST SPRING Grant Number JPMJSP2140.